# Multi-scale phylodynamic modelling of rapid punctuated pathogen evolution

### v2

@@ D.2.3 Incorporating compound immunity in infection probability
Here, we incorporate the compound immunity, against infection ($M^{\theta}$) and forward transmission ($M^{\iota}$), within the infection probability, extending the NPI-affected probability of infection (Equation 11 in Section C.2).
@@
The probability that susceptible agent $i$ becomes ill (i.e., symptomatic), originally defined in Equation 10, is then updated as:
@@ Appendix E Parametrisation and additional results
Figures 19 to 22 illustrate simulated epidemic dynamics, as well as the dynamics of transmissibility and mutation accumulation across populations of varying sizes: small (1.7 million), medium (8.0 million), and large (25.4 million). These figures provide insights into population health states, infection and immunisation histories, and viral transmissibility trends, aligned with phylogenetic dynamics observed in different realisations across diverse population scenarios.
Figures 23 to 25 present pairwise Hamming distance dynamics generated using identical phylogenetic parameters across three different population sets. Fig. 23 shows alternative evolution dynamics relative to those shown in Fig. 9 (A, brown dashed line). The high variability of genomic diversity dynamics across different realisations results from the stochastic emergence of variants; hence, we refrain from averaging genomic diversity dynamics across realisations. Nonetheless, the pattern of genomic diversity dynamics aligns well with the trends discussed in Section 2.2, which includes periods of drift, rapid rise, and abrupt collapse. To assess the stationarity of genomic diversity dynamics, we performed the Augmented Dickey–Fuller test. Fig. 27 shows that the p-value for stationarity associated with genomic diversity dynamics varies significantly with population size, highlighting the increased difficulty for variants to spread and become dominant in larger populations.
Figure 20: Simulated dynamics of the average transmissibility, i.e., fitness ($K$, olive bars, y-axis on the left), and the accumulated mutations ($\widehat{D}$, solid brown line, y-axis on the right) for a 1.7 million population from 2020 to 2026. Panels (A)-(D) are profiles plotted using four different realisations.
Figure 21: Simulated dynamics of the average transmissibility, i.e., fitness ($K$, olive bars, y-axis on the left), and the accumulated mutations ($\widehat{D}$, solid brown line, y-axis on the right) for an 8.0 million population from 2020 to 2026. Panels (A)-(D) are profiles plotted using four different realisations.
Figure 22: Simulated dynamics of the average transmissibility, i.e., fitness ($K$, olive bars, y-axis on the left), and the accumulated mutations ($\widehat{D}$, solid brown line, y-axis on the right) for a 25.4 million population from 2020 to 2026. Panels (A)-(D) are profiles plotted using four different realisations.
@@
Figure 25: Simulated dynamics of the average pairwise Hamming distance between two randomly selected genomes from infected hosts in a population of 25.4 million. Approximately 10,000 pairs of genomes are randomly sampled at each simulation time point from 2020 to 2026 (Capability 2(i)). Opaque green lines represent the ensemble of all realisations and the solid green line shows the dynamics of one realisation only.
@@ F.1 Chronic infections
Fig. 26 differentiates between the COVID-19 pandemic simulation scenarios with and without chronic infections. Notably, the absence of chronic infections results in a reduced number of recurrent incidence waves (Fig. 26 (A)) and significantly reduced fluctuations in the genomic diversity $\overline{D}$, producing lower variability (i.e., a smaller standard deviation) across realisations (Fig. 26 (D)). Furthermore, the growth of both transmissibility (i.e., fitness $K$) and the accumulated mutations $\widehat{D}$ is significantly slower (Fig. 26 (B) and (C)), impeding the emergence of high-fitness variants.
We also explored whether the pathogen phylodynamics are correlated with (1) fractions of chronic infection, and (2) strength of the within-host selective pressure in chronically infected hosts. Fig. 29 shows that a hundred-fold increase in the chronic infection fraction (e.g., from 0.05% to 5% of the population) yields only negligible impact on the incidence curves and phylodynamics. This indicates a ceiling effect of the chronic infection fraction on the incidence, fitness, and accumulated mutations. In contrast, Fig. 30 presents a clear positive correlation between the within-host selective pressure and the phylodynamic characteristics. That is, a higher selective pressure (i.e., selecting top 10% of mutated genome candidates out of their ranked list, $X=10$, $M=100$) leads to higher incidence peaks and a more rapid increase in the pathogen fitness and accumulated mutations.
@@ F.2 Stationarity of genomic diversity
The genomic diversity traced in empirical data from 2020 to 2024 shows fluctuating yet relatively stationary dynamics, without notable trends of increasing or decreasing diversity, as illustrated in Fig. 2 (C)). The Augmented Dicky-Fuller (ADF) test (detailed in Section 4) applied for the empirical genomic diversity produced p-value of 0.024, indicating stationary at the significance level of 0.05.
However, stationarity of the simulated genomic diversity is harder to establish for larger population sizes, as illustrated in Fig. 27. For a small population of 1.7 million, the genomic diversity dynamics are closer to stationary (p-value of 0.078). This outcome is robust to variations of the fitness contributions weight table, as well as changes in the within-host selective pressure, as shown in Fig. 28. For larger population sizes, the ADF tests produced larger p-values of 0.317 (8 million) and 0.730 (25.4 million), indicating a progressive loss of stationarity.
Figure 26: Simulated dynamics in scenarios with the fraction of agents susceptible to chronic infections being 0.1% of the entire population (blue curves), and without chronic infections (red curves). The profiles trace detected (A) incidence, (B) average pathogen transmissibility, interpreted as fitness $K$, (C) accumulated mutations ($\widehat{D}$), and (D) genomic diversity ($\overline{D}$) in the population of 1.7 million, over six simulation years. The shaded area represents the range of one standard deviation from the mean values. In (A) to (C), the solid lines denote the mean value across 30-50 realisations. In (D), the dashed line denotes the pairwise Hamming distance from one realisation only, given the high variability between realisations.
Figure 27: Statistical estimation of stationarity of the time evolution of the average pairwise Hamming distance between two randomly selected genomes. Approximately 10,000 pairs of genome sequences are randomly sampled at each simulation time point from 2020 to 2026 for three simulation scenarios with different population sizes of 1.7 million (A), 8 million (B), and 25.4 million (C) (Capability 3(ii)).
Figure 28: Statistical estimation of the stationarity of the average pairwise Hamming distance between two randomly selected genomes simulated with populations of 1.7 million (A) and 8.0 million (B). Approximately 10,000 pairs of genome sequences are randomly sampled at each simulation time point from 2020 to 2026 for four simulation scenarios by varying amino acid weight table and/or within-host selective pressure (see bar legends). We tested two different fitness weight tables following two different normal distributions: (i) BT: the baseline setting, where each amino acid contribution in spike and non-spike codon positions is sampled from the normal distributions N(0, 0.085) and N(0, 0.07), respectively, and (ii) AT: the alternative setting, where each amino acid contribution in spike and non-spike codon positions is sampled from the normal distributions N(0, 0.0489) and N(0, 0.0454), respectively. Within-host selective pressure is varied for both typical infected hosts and chronically infected hosts, using notation separated by a vertical bar (e.g., $Top99|Top30$ represents a setting where selective pressure at top 99 for typically infected hosts (X=99, M=100) and at top 30 for chronically infected hosts (X=30, M=100).
@@ Appendix G Sensitivity analysis
We performed sensitivity analysis by tracing the output variables of interest in response to changes in one input variable while keeping the other inputs specified at default values. In this study, we have four output variables of interest: incidence, transmissibility fitness, accumulated mutations, and genomic diversity. Figures 29 and 30 show simulated dynamics of the output variables of interest by varying two input parameters: the fraction of chronically infected hosts and the within-host selective pressure in chronically infected hosts, respectively.
Figure 29: Simulated dynamics of the detected incidence (I-A), average pathogen fitness (I-B), average accumulated mutations $\widehat{D}$ (II-A), and average genomic diversity $\overline{D}$ (II-B) in a population of 1.7 million. These measures are traced while the fraction of individuals susceptible to chronic infection, with strong positive within-host selective pressure, is varied from 0.0 to 0.05 of the total population.
Figure 30: Simulated dynamics of (A) detected incidence, (B) average pathogen fitness, and (C) average accumulated mutations $\widehat{D}$ in a population of 1.7 million. The within-host selective pressure for chronically infected hosts, starting from day 60 of infection, is varied by selecting, in each simulation cycle, the mutated genomes from the top 10% (high selectivity, $X=10$, $M=100$) to the top 90% (low selectivity, $X=90$, $M=100$) of the ranked list of mutated genome candidates.
@@ Appendix H Computational complexity and implementation
The multi-scale phylodynamic simulator, PhASE TraCE, builds upon our open-source agent-based epidemiological simulator, AMTRaC-19 [57], written in C++. The architecture of PhASE TraCE includes four layers (Fig. 12), with the immunological and phylogenetic layers being novel additions, which significantly extend the capabilities of AMTRaC-19. PhASE TraCE is designed to simulate the dynamic interaction between pathogen transmission and evolution over a prolonged timeframe over multiple years (unlike AMTRaC-19 which was typically used to simulate 6-9 months of a pandemic). The new modelling capabilities and simulation requirements increase the computational complexity, presenting four specific computational challenges:
@@ H.1 Performance and scalability
To improve computational efficiency, PhASE TraCE utilises multi-threading processing provided by the C++ OpenMP library. Figure 31 traces the average processing time of simulation runs performed on a high-performance computing cluster, for three different population sizes, ranging from approximately 230,000 agents (Figure 31.A), to approximately 1,700,000 agents (Figure 31.B), and approximately 8,000,000 agents (Figure 31.C).
The average processing time per simulation day increases during the simulation: this occurs due to the increased memory (see the top row of Figure 31). Nevertheless, there is a consistent reduction of the processing time when more CPUs are used per job. We note that, for a given number of CPUs, the average cumulative processing time linearly increases in proportion to the population size. For example, Figure 32 shows that when 8 CPUs are used, a simulation of a 365-day period across a population with 8 million agents completes in approximately 20,000 seconds (over 5.5 hours).
@@ H.2 Memory utilisation
An increasing size of the artificial population requires higher computation resources. This is because the stochastically generated agents are assigned several static demographic attributes, which are preserved during the simulation, and a number of dynamic phylogenetic and immunological attributes, which change during the simulation. These attributes contribute to both fixed memory and dynamic memory utilisation.
@@
As the simulation progresses, the number of stored immunological records grows significantly due to the increased number of infections and administered vaccinations. Figure 33 shows how the dynamic memory usage (in Gigabytes) increases with the larger number of stored immunological records across the population of 8 million agents.

### v3

@@ D.2.3 Incorporating compound immunity in infection probability
Here, we incorporate the compound immunity, against infection ($M^{\theta}$) and forward transmission ($M^{\iota}$), within the infection probability, extending the NPI-affected probability of infection (Eq 11 in Section C.2).
@@
The probability that susceptible agent $i$ becomes ill (i.e., symptomatic), originally defined in Eq 10, is then updated as:
@@ Appendix E Parametrisation and additional results
Figure 19: Detected incidence: log scale. (A) Simulated incidence on a log scale in population sets of 1.7 million (red), 8 million (blue), and 25.4 million (green). This figure corresponds to Fig 6A on a linear scale. (B) Worldwide detected incidence [14], measured as new weekly cases per million (solid black line) and new daily cases per million (dashed black line).
@@
Figs 20 to 23 illustrate simulated epidemic dynamics, as well as the dynamics of transmissibility and mutation accumulation across populations of varying sizes: small (1.7 million), medium (8.0 million), and large (25.4 million). These figures provide insights into population health states, infection and immunisation histories, and viral transmissibility trends, aligned with phylogenetic dynamics observed in different realisations across diverse population scenarios.
Figs 25 to 27 present pairwise Hamming distance dynamics generated using identical phylogenetic parameters across three different population sets. Fig 25 shows alternative evolution dynamics relative to those shown in Fig. 9 (A, brown dashed line). The high variability of genomic diversity dynamics across different realisations results from the stochastic emergence of variants; hence, we refrain from averaging genomic diversity dynamics across realisations. Nonetheless, the pattern of genomic diversity dynamics aligns well with the trends discussed in Section 2.2, which includes periods of drift, rapid rise, and abrupt collapse. To assess the stationarity of genomic diversity dynamics, we performed the Augmented Dickey–Fuller test. Fig 29 shows that the p-value for stationarity associated with genomic diversity dynamics varies significantly with population size, highlighting the increased difficulty for variants to spread and become dominant in larger populations.
Figure 21: Simulated dynamics of the average transmissibility for a 1.7 million population. Fitness ($K$, olive bars, y-axis on the left), and the accumulated mutations ($\widehat{D}$, solid brown line, y-axis on the right) are plotted from 2020 to 2026. Panels (A)-(D) are profiles plotted using four different realisations.
Figure 22: Simulated dynamics of the average transmissibility for an 8.0 million population. Fitness ($K$, olive bars, y-axis on the left), and the accumulated mutations ($\widehat{D}$, solid brown line, y-axis on the right) are plotted from 2020 to 2026. Panels (A)-(D) are profiles plotted using four different realisations.
Figure 23: Simulated dynamics of the average transmissibility for a 25.4 million population. Fitness ($K$, olive bars, y-axis on the left), and the accumulated mutations ($\widehat{D}$, solid brown line, y-axis on the right) are plotted from 2020 to 2026. Panels (A)-(D) are profiles plotted using four different realisations.
Figure 24: Simulated dynamics of the average transmissibility across different genome regions for a 1.7 million population. (A) Fitness $K$ per codon, and (B) accumulated mutations $\widehat{D}$ per codon. Five genome regions are plotted: overall genome (red), spike and non-epitope region (blue), spike and epitope (green), non-spike and epitope (purple), and non-spike and non-epitope (pink).
@@
Figure 27: Simulated dynamics of the average pairwise Hamming distance between two randomly selected genomes from infected hosts in a population of 25.4 million. Approximately 10,000 pairs of genomes are randomly sampled at each simulation time point from 2020 to 2026 (Capability 2(i)). Opaque blue lines represent the ensemble of all realisations and the solid blue line shows the dynamics of one realisation only.
@@ F.1 Chronic infections
Fig 28 differentiates between the COVID-19 pandemic simulation scenarios with and without chronic infections. Notably, the absence of chronic infections results in a reduced number of recurrent incidence waves (Fig 28 (A)) and significantly reduced fluctuations in the genomic diversity $\overline{D}$, producing lower variability (i.e., a smaller standard deviation) across realisations (Fig  28 (D)). Furthermore, the growth of both transmissibility (i.e., fitness $K$) and the accumulated mutations $\widehat{D}$ is significantly slower (Fig  28 (B) and (C)), impeding the emergence of high-fitness variants.
We also explored whether the pathogen phylodynamics are correlated with (1) fractions of chronic infection, and (2) strength of the within-host selective pressure in chronically infected hosts. Fig 31 shows that a hundred-fold increase in the chronic infection fraction (e.g., from 0.05% to 5% of the population) yields only negligible impact on the incidence curves and phylodynamics. This indicates a ceiling effect of the chronic infection fraction on the incidence, fitness, and accumulated mutations. In contrast, Fig 32 presents a clear positive correlation between the within-host selective pressure and the phylodynamic characteristics. That is, a higher selective pressure (i.e., selecting top 10% of mutated genome candidates out of their ranked list, $X=10$, $M=100$) leads to higher incidence peaks and a more rapid increase in the pathogen fitness and accumulated mutations.
@@ F.2 Stationarity of genomic diversity
The genomic diversity traced in empirical data from 2020 to 2024 shows fluctuating yet relatively stationary dynamics, without notable trends of increasing or decreasing diversity, as illustrated in Fig 2 (C)). The Augmented Dicky-Fuller (ADF) test (detailed in Section 4) applied for the empirical genomic diversity produced a p-value of 0.024, indicating stationarity at the significance level of 0.05.
However, stationarity of the simulated genomic diversity is harder to establish for larger population sizes, as illustrated in Fig 29. For a small population of 1.7 million, the genomic diversity dynamics are closer to stationary (p-value of 0.078). This outcome is robust to variations of the fitness contributions weight table, as well as changes in the within-host selective pressure, as shown in Fig 30. For larger population sizes, the ADF tests produced larger p-values of 0.317 (8 million) and 0.730 (25.4 million), indicating a progressive loss of stationarity.
@@ Appendix G Sensitivity analysis
We performed sensitivity analysis by tracing the output variables of interest in response to changes in one input variable while keeping the other inputs specified at default values. In this study, we have four output variables of interest: incidence, transmissibility fitness, accumulated mutations, and genomic diversity. Figs 31 and 32 show simulated dynamics of the output variables of interest by varying two input parameters: the fraction of chronically infected hosts and the within-host selective pressure in chronically infected hosts, respectively.
@@ Appendix H Computational complexity and implementation
The multi-scale phylodynamic simulator, PhASE TraCE, builds upon our open-source agent-based epidemiological simulator, AMTRaC-19 [59], written in C++. The architecture of PhASE TraCE includes four layers (Fig 12), with the immunological and phylogenetic layers being novel additions, which significantly extend the capabilities of AMTRaC-19. PhASE TraCE is designed to simulate the dynamic interaction between pathogen transmission and evolution over a prolonged timeframe over multiple years (unlike AMTRaC-19 which was typically used to simulate 6-9 months of a pandemic). The new modelling capabilities and simulation requirements increase the computational complexity, presenting four specific computational challenges:
@@ H.1 Performance and scalability
To improve computational efficiency, PhASE TraCE utilises multi-threading processing provided by the C++ OpenMP library. Fig 33 traces the average processing time of simulation runs performed on a high-performance computing cluster, for three different population sizes, ranging from approximately 230,000 agents (Fig  33.A), to approximately 1,700,000 agents (Fig  33.B), and approximately 8,000,000 agents (Fig  33.C).
The average processing time per simulation day increases during the simulation: this occurs due to the increased memory (see the top row of Fig  33). Nevertheless, there is a consistent reduction of the processing time when more CPUs are used per job. We note that, for a given number of CPUs, the average cumulative processing time linearly increases in proportion to the population size. For example, Fig 34 shows that when 8 CPUs are used, a simulation of a 365-day period across a population with 8 million agents completes in approximately 20,000 seconds (over 5.5 hours).
@@ H.2 Memory utilisation
An increasing size of the artificial population requires higher computational resources. This is because the stochastically generated agents are assigned several static demographic attributes, which are preserved during the simulation, and a number of dynamic phylogenetic and immunological attributes, which change during the simulation. These attributes contribute to both fixed memory and dynamic memory utilisation.
@@
As the simulation progresses, the number of stored immunological records grows significantly due to the increased number of infections and administered vaccinations. Fig 35 shows how the dynamic memory usage (in Gigabytes) increases with the larger number of stored immunological records across the population of 8 million agents.
Figure 28: Counterfactual scenario: chronic infections. Simulated dynamics in scenarios with the fraction of agents susceptible to chronic infections being 0.1% of the entire population (blue curves), and without chronic infections (red curves). The profiles trace detected (A) incidence, (B) average pathogen transmissibility, interpreted as fitness $K$, (C) accumulated mutations ($\widehat{D}$), and (D) genomic diversity ($\overline{D}$) in the population of 1.7 million, over six simulation years. The shaded area represents the range of one standard deviation from the mean values. In (A) to (C), the solid lines denote the mean value across 30-50 realisations. In (D), the dashed line denotes the pairwise Hamming distance from one realisation only, given the high variability between realisations.
Figure 29: Statistical estimation of stationarity. The stationarity of the time evolution of the average pairwise Hamming distance is computed between two randomly selected genomes. Approximately 10,000 pairs of genome sequences are randomly sampled at each simulation time point from 2020 to 2026 for three simulation scenarios with different population sizes of 1.7 million (A), 8 million (B), and 25.4 million (C) (Capability 3(ii)).
Figure 30: Statistical estimation of the stationarity of the average pairwise Hamming distance between two randomly selected genomes simulated with with different population sizes. (A) populations of 1.7 million and (B) 8.0 million. Approximately 10,000 pairs of genome sequences are randomly sampled at each simulation time point from 2020 to 2026 for four simulation scenarios by varying amino acid weight table and/or within-host selective pressure (see bar legends). We tested two different fitness weight tables following different normal distributions: (i) BT: the baseline setting, where each amino acid contribution in spike and non-spike codon positions is sampled from the normal distributions N(0, 0.085) and N(0, 0.07), respectively, and (ii) AT: the alternative setting, where each amino acid contribution in spike and non-spike codon positions is sampled from the normal distributions N(0, 0.0489) and N(0, 0.0454), respectively. Within-host selective pressure is varied for both typical infected hosts and chronically infected hosts, using notation separated by a vertical bar (e.g., $Top99|Top30$ represents a setting where selective pressure at top 99 for typically infected hosts (X=99, M=100) and at top 30 for chronically infected hosts (X=30, M=100).
Figure 31: Sensitivity analysis: the fractions of chronic infections. Simulated dynamics of the detected incidence (I-A), average pathogen fitness (I-B), average accumulated mutations $\widehat{D}$ (II-A), and average genomic diversity $\overline{D}$ (II-B) (one realisation only), in a population of 1.7 million. These measures are traced while the fraction of individuals susceptible to chronic infection, with strong positive within-host selective pressure, is varied from 0.0 to 0.05 of the total population.
Figure 32: Sensitivity analysis: within-host selective pressure. Simulated dynamics of (A) detected incidence, (B) average pathogen fitness, and (C) average accumulated mutations $\widehat{D}$ in a population of 1.7 million. The within-host selective pressure for chronically infected hosts, starting from day 60 of infection, is varied by selecting, in each simulation cycle, the mutated genomes from the top 10% (high selectivity, $X=10$, $M=100$) to the top 90% (low selectivity, $X=90$, $M=100$) of the ranked list of mutated genome candidates.